1))) According to Newton’s laws, an object will remain stationary or move at constant velocity unless acted

by a force. If a force (F) is applied, the object will accelerate according to F=m·a, where “m” is the mass

of the object and “a” is its acceleration. The unit of force is the Newton (N), N is kg·m/sec

a) Consider a stationary object with a mass of 5 Kg, acted on by a force of 10 N. What is its acceleration?

b) If that acceleration continues, how long would it take for the object to reach a speed of 40 m/sec?
Answer #1

(a)F=10N

mass(m)=5kg

Ignoring the friction between object and surface.

According to Newton's law-

F=ma

Where a is acceleration.

10=5a

a=2meter /sec​​​​2​​​ Ans.

(b)object starts moving from stationary,so initial velocity-

u=0

Final velocity -

v=40m/sec.

Using -

v=u+at..............(1)

a=2m/sec​​​​2  (from first question)

putting all value in equation (1)-

40=0+2t

t=40/2=20sec Ans.